# Experimental Observation of Motion of Ions in a Resonantly Driven Plasma Wakefield Accelerator

### v1

@@ Abstract
We also note that the effect of the motion of plasma ions extends beyond that on bunch emittance and energy gain. It potentially imposes a limitation on the repetition rate of the acceleration process due to the time it takes for the energy of wakefields to dissipate Zgadzaj2020 and correspondingly, the time the plasma takes to recover from the excitation-acceleration process that leaves energy in the plasma. This recovery time includes reaching again a uniform plasma density, i.e., uniform electron, ion, and neutral densities, as measured in Ref. D’Arcy2022 .
@@
The combination of experimental and simulation results presented in this Letter clearly show the effect of the motion of ions on the development of self-modulation, in line with expectations. This effect (formation of a bunch tail) is caused by the ponderomotive force of the wakefields. The experimental results confirm that the effect depends inversely on the mass of the ions (helium, argon, xenon), and depends directly on the amplitude of the wakefields (for a fixed ion mass). Simulation results show that the effect of motion of ions occurs later along the bunch than where wakefields reach their maximum amplitude. This is essential for future acceleration experiments at AWAKE aimed at reaching large electron bunch energy gain ($\cong$50 GeV). The dependence observed with mass, and that with amplitude, are in agreement with those of theoretical and simulation models PhysRevLett.109.145005 ; 10.1063/1.4876620 . These will be used to evaluate the possible effect of the motion of ions, e.g. in single and multiple drivers wakefield accelerators that may use light ions to avoid multiple ionization levels in the strong fields of the intense driver and witness beams.

### v3

@@ Abstract
We also note that the long-timescale ($>1$ ns) effect of the motion of plasma ions extends beyond that on bunch emittance and energy gain. It potentially imposes a limitation on the repetition rate of the acceleration process due to the time it takes for the energy of wakefields to dissipate Zgadzaj2020 and correspondingly, the time the plasma takes to recover from the excitation-acceleration process that leaves energy in the plasma. This recovery time includes reaching again a uniform plasma density, i.e., uniform electron, ion, and neutral densities, as measured in Ref. D’Arcy2022 .
@@
The combination of experimental and simulation results presented in this Letter clearly show an effect of the motion of ions within a single wakefield event and for the first time with different ion masses. This effect (formation of a bunch tail) is caused by the ponderomotive force of the wakefields. The experimental results confirm that the effect depends inversely on the mass of the ions (helium, argon, xenon), i.e. appears first with lighter ions, and increases with the amplitude of the wakefields. The dependence observed with mass, and that with amplitude, are in agreement with those of theoretical and simulation models PhysRevLett.109.145005 ; 10.1063/1.4876620 ; lotovion . The scaling with $m_{i}^{-1/3}$ can be used to evaluate the possible effect of the motion of ions, e.g. in single and multiple drivers wakefield accelerators that may use light ions to avoid multiple ionization levels in the strong fields of the intense driver and witness beams.